# Searching for local associations while controlling the false discovery rate

### v1

@@ 3.3 Data Analysis: Genetic Underpinning of WHR Across Sexes
We study the genetic basis of Waist-Hip Ratio (WHR) across sexes using the UK Biobank data (Bycroft et al., 2018); see Appendix A5 for details. Previous studies have examined sex-specific effects for WHR or WHR adjusted for BMI, identifying between 7 and 44 sex-specific loci (Heid et al., 2010; Randall et al., 2013; Winkler et al., 2015; Shungin et al., 2015; Funkhouser et al., 2020). Given the relatively small number of loci discovered in prior research, we anticipate a high proportion of global effects.
We work with $n=346,119$ unrelated British individuals for whom we have genotypes ($p=591,513$ variants) and measures on sex (46.35% male), age, height, weight, BMI, waist and hip circumference. We define WHR = waist circumference / hip circumference.
We apply aLKF using the specialized implementation described in Appendix A2.3, designed for large-scale GWAS data with a very large number $p$ of explanatory variables. This implementation leverages lasso models and two key computational shortcuts: pre-screening and batch computation of importance scores. The lasso models adjust for the following covariates: age, age squared, the first five genetic principal components, and sex (if applicable). The target FDR level is $\alpha=0.1$. Similar to the simulation in Section 3.2 and as discussed in Appendix A2.3.4 and A5, inference is conducted using blocks of SNPs, with median width equal to 81 kilobases (kb), as the units of inference.
To learn an informative partition function, after pre-screening, we fit a lasso model including all possible pairwise interactions between each of the pre-screened SNPs and sex; see Appendix A2.3.4 for further details on the process. This leads to the definition of 15,001 block-level hypotheses for the full sample and 175 hypotheses each for males and females.
@@ 4 Discussion
The data analysis and simulations presented in this paper all reflect the results of one knockoff generation, with associated randomness. In practical applications, one might want to resort to multiple knockoff to increase the stability of selection (Ren et al., 2023).
The first step of our procedure consists in identifying the relevant “neighborhoods” in which to test for association between $Y$ and each of the $X_{j}$. In the examples considered in this paper, we found that the lasso with interactions can be useful in defining a partition $\nu_{j}(Z)$. We also noted that different penalization levels for main effects and interactions might be appropriate to navigate the trade-off between the precision of the hypotheses we decided to test with the power to reject them, linked to the size of the subgroup defined by $\nu_{j}(Z)$ (see Appendix A2.3.1). A better balance between these might be achieved with a resolution-adaptive procedure as the one described in Gablenz and Sabatti (2024).
@@ Acknowledgments
Recall the high-level ideas of the procedure described in Section 2.5, also visualized schematically in Figure A2. For each pair $(j,\ell)$, we extract the scores $(T_{j,\ell},\tilde{T}_{j,\ell})$ from a local predictive model $\hat{\psi}^{(j,\ell)}$ trained using information from the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and the local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, which includes only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$. We now outline this implementation in more detail.
Figure A2: Schematic of the Adaptive Local Knockoff Filter. The relevant variables and associated statistics for different local environments are shown in different colors. The original variables are shown in darker shades and the knockoffs in lighter shades.
@@ 11                   For each $l\in[p]$, define the weight $\smash{\lambda_{l}=\lambda(1-\xi)+\xi\pi_{l}}$;
Figure A3: Schematic of Adaptive Local Knockoff Filter, focusing on the computation of the statistics for the first (blue) and fifth (green) local environments. The test statistics are based on a simple empirical prior for the two variables that is learnt from the cloaked data set with randomly swapped knockoffs. Other details are as in Figure A2.
@@ 2 $Z_{3}=1$ $X_{2}$ influences $Y$ if $Z_{3}=1$
Algorithms 4 and 5 summarize this specialized implementation of our method, with additional details provided in Appendices A2.3.1A2.3.3. Furthermore, Appendix A2.3.4 describes how our method can be applied after partitioning the genome into blocks of SNPs. This accounts for strong linkage disequilibrium (LD) among nearby variants on the same chromosome, enabling a multi-resolution analysis (Sesia et al., 2020, 2021).
@@ 2 $Z_{1}=1$
To address this complication in GWAS data, Sesia et al. (2020) and Sesia et al. (2021) applied the global knockoff filter of Candès et al. (2018) at different levels of resolution. Specifically, they partitioned the genome into contiguous blocks of variants in strong linkage disequilibrium and tested for conditional independence with the phenotype at the block level, using these blocks as the units of inference rather than individual SNPs.
The same strategy can be seamlessly integrated with our aLKF method.
Starting with a pre-defined partition $\mathcal{B}$ of the genome into contiguous blocks of SNPs $b\in\mathcal{B}$ at the desired resolution, along with corresponding knockoff genotypes, as described in Sesia et al. (2020) and Sesia et al. (2021), we can apply our method as outlined so far, with four key modifications:
@@
    The data cloaking step using the random matrix $\mathbf{V}$ is performed at the block level, assigning the same values of $V_{ij}$ to all SNPs $j$ within a block $b\in\mathcal{B}$. This reflects the idea that the units of inference are now blocks of SNPs and not single SNPs.
@@
    Each block of SNPs $b$ is assigned a single partition function $\hat{\nu}_{b}$, rather than having distinct partition functions for individual SNPs.
@@
    While pre-screening nominally selects a subset of individual variants, since inference is then performed at the block level, the batch-local predictive models are fitted using not only the pre-screened variants but also all other variants in the same block.
@@
    For each combination of $b$ and $\ell\in\hat{L}_{b}$, a single pair of local scores $(T_{b,\ell},\tilde{T}_{b,\ell})$ is computed by summing the single-SNP scores corresponding to the individual SNPs within that block. Then, the corresponding test statistic is $W_{b,\ell}=T_{b,\ell}-\tilde{T}_{b,\ell}$.
Finally, the discoveries obtained with our method will be interpreted as highlighting blocks of SNPs displaying significant associations with the outcome within certain subgroups determined by specific covariate values.
@@
Background on block-level inference.
As discussed in Appendix A2.3.4 and A2.3.4, Sesia et al. (2021) partition the individual SNPs into contiguous groups at seven levels of resolution based on complete linkage hierarchical clustering. The group sizes range from single SNPs to 425-kilobase (kb) wide groups. For the real data analysis we choose the resolution with median group width 81 kb and for the simulation we choose a resolution with median group width 3 kb. In the simulation, for simplicity, we only choose a single variant within each of these groups to be nonzero. Since inference is based on the block level, the presence of a single nonzero SNP within a block renders the conditional independence hypothesis to be false at the block level as well.
@@
After the construction of knockoff variables and the assembly of cloaked data, the entire computational pipeline for the real data application in Section 3.3 can be run within approximately 15 hours. All analyses are conducted in R/4.2.0 and use the “bigstatsr” and “bigsnpr” packages (Privé et al., 2018) for handling genetic data.
@@
Figure A6: Same general simulation setting as A4, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
Figure A7: Same simulation setting as  A6 with different levels of sparsity.
@@
Figure A10: Same simulation setting as in A8, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.

### v2

@@ 3.3 Data Analysis: Genetic Underpinning of WHR Across Sexes
We study the genetic basis of Waist-Hip Ratio (WHR) across sexes using the UK Biobank data (Bycroft et al., 2018). Previous studies have examined sex-specific effects for WHR or WHR adjusted for BMI, identifying between 7 and 44 sex-specific loci (Heid et al., 2010; Randall et al., 2013; Winkler et al., 2015; Shungin et al., 2015; Funkhouser et al., 2020). Given the relatively small number of loci discovered in prior research, we anticipate a high proportion of global effects.
We work with $n=346,119$ unrelated British individuals for whom we have genotypes ($p=591,513$ variants) and measures on sex (46.35% male), age, height, weight, BMI, waist and hip circumference. We define WHR = waist circumference / hip circumference. After the construction of knockoff variables with the approach of Sesia et al. (2021) and the assembly of cloaked data, the analysis described below takes approximately 15 hours on a computing cluster. We refer to Appendix A5 for additional computational details.
We apply aLKF using the specialized implementation described in Appendix A2.3, designed for large-scale GWAS data with a very large number $p$ of explanatory variables. This implementation leverages lasso models and two key computational shortcuts: pre-screening and batch computation of importance scores. The lasso models adjust for the following covariates: age, age squared, the first five genetic principal components, and sex (if applicable). The target FDR level is $\alpha=0.1$. As in the simulation in Section 3.2, and detailed in Appendices A2.3.4 and A5, inference is performed at the level of SNP groups, with a median group width of 81 kilobases (kb). In real data, this resolution offers a practical balance between power and precision of the tested hypotheses (Sesia et al., 2020, 2021; Gablenz and Sabatti, 2025).
To learn an informative partition function, after pre-screening, we fit a lasso model including all possible pairwise interactions between the pre-screened SNPs and sex; see Appendix A2.3.4 for further details. This leads to the definition of 15,001 SNP-group-level hypotheses for the full sample and 175 hypotheses each for males and females.
@@ 4 Discussion
The data analysis and simulations presented in this paper all reflect the results of one knockoff generation, with associated randomness. In practice, one may wish to use multiple knockoff realizations to improve stability (He et al., 2021). An interesting direction for future research would be to investigate how the de-randomization approach of Ren et al. (2023), which leverages e-values (Wang and Ramdas, 2022), could be adapted to our setting.
The first step of our procedure consists in identifying the relevant “neighborhoods” in which to test for association between $Y$ and each of the $X_{j}$. In the examples considered in this paper, we found that the lasso with interactions can be useful in defining a partition $\nu_{j}(Z)$. We also noted that different penalization levels for main effects and interactions might be appropriate to navigate the trade-off between the precision of the tested hypotheses with the power to reject them, linked to the size of the subgroup defined by $\nu_{j}(Z)$ (see Appendix A2.3.1). A better balance between these might be achieved with a resolution-adaptive procedure as the one described in Gablenz and Sabatti (2025), leveraging e-values.
@@ Acknowledgments
Recall the high-level ideas of the procedure described in Section 2.5, also visualized schematically in Figure 1. For each pair $(j,\ell)$, we extract the scores $(T_{j,\ell},\tilde{T}_{j,\ell})$ from a local predictive model $\hat{\psi}^{(j,\ell)}$ trained using information from the cloaked dataset $\tilde{\mathcal{D}}(\mathbf{V})$ and the local dataset $\mathcal{D}^{(j,\ell)}=(\mathbf{X}_{j},\tilde{\mathbf{X}}_{j},\mathbf{Y},\mathbf{Z})^{(j,\ell)}$, which includes only observations $(X_{j},\tilde{X}_{j},Y,Z)$ for individuals $i\in[n]$ with $\nu_{j}(Z^{i})=\ell$. We now outline this implementation in more detail.
@@ 11                   For each $l\in[p]$, define the weight $\smash{\lambda_{l}=\lambda(1-\xi)+\xi\pi_{l}}$;
Figure A2: Schematic of Adaptive Local Knockoff Filter, focusing on the computation of the statistics for the first (blue) and fifth (green) local environments. The test statistics are based on a simple empirical prior for the two variables that is learned from the cloaked data set with randomly swapped knockoffs. Other details are as in Figure 1.
@@ 2 $Z_{3}=1$ $X_{2}$ influences $Y$ if $Z_{3}=1$
Algorithms 4 and 5 summarize this specialized implementation of our method, with additional details provided in Appendices A2.3.1A2.3.3. Furthermore, Appendix A2.3.4 describes how our method can be applied after partitioning the genome into groups of SNPs. This accounts for strong linkage disequilibrium (LD) among nearby variants on the same chromosome, enabling a multi-resolution analysis (Sesia et al., 2020, 2021).
@@ 2 $Z_{1}=1$
To address this complication in GWAS data, Sesia et al. (2020) and Sesia et al. (2021) applied the global knockoff filter of Candès et al. (2018) at different levels of resolution. Specifically, they partitioned the genome into contiguous groups of variants in strong linkage disequilibrium and tested for conditional independence with the phenotype at the level of groups of SNPs. These groups of SNPs are then used as the units of inference rather than individual SNPs.
Specifically, let $\mathcal{B}$ be a partition of the genome into contiguous groups of SNPs $b\in\mathcal{B}$. The equivalent environment-level hypothesis at the group level, see Equation (6), is
¯$\bar{\mathcal{H}}_{b,\ell}(\nu):Y^{(b,\ell)}\perp\!\!\!\perp X^{(b,\ell)}_{b}\mid X^{(b,\ell)}_{-b},Z^{(b,\ell)}\;\;\;\;b\in\mathcal{B},\ell\in[L_{b}].$
This hypothesis is analogous to the one in Equation (6), except that inference is at the level of groups of SNPS, rather than single SNPs.
Similarly, the exchangeability is required to hold for groups of SNPs, see also Equation (7):
[…]
Sesia et al. (2021) construct valid knockoffs for groups of SNPs using hidden Markov models, see Sesia et al. (2021) for further details. They construct groups of SNPs using complete-linkage hierarchical clustering. We can seamlessly integrate the genetic group level strategy with our aLKF method. To test the hypothesis described above, starting with a pre-defined partition $\mathcal{B}$ of the genome into contiguous groups of SNPs $b\in\mathcal{B}$ at the desired resolution, along with corresponding knockoff genotypes, as described in Sesia et al. (2020) and Sesia et al. (2021), we can apply our method as outlined so far, with four key modifications:
@@
    The data cloaking step using the random matrix $\mathbf{V}$ is performed at the level of groups of SNPs, assigning the same values of $V_{ij}$ to all SNPs $j$ within a group of SNPs $b\in\mathcal{B}$. This reflects the idea that the units of inference are now groups of SNPs and not single SNPs.
@@
    Each group of SNPs $b$ is assigned a single partition function $\hat{\nu}_{b}$, rather than having distinct partition functions for individual SNPs.
@@
    While pre-screening nominally selects a subset of individual variants, since inference is then performed at the level of groups of SNPs, the batch-local predictive models are fitted using not only the pre-screened variants but also all other variants in the same genetic group.
@@
    For each combination of $b$ and $\ell\in\hat{L}_{b}$, a single pair of local scores $(T_{b,\ell},\tilde{T}_{b,\ell})$ is computed by summing the single-SNP scores corresponding to the individual SNPs within that group. Then, the corresponding test statistic is $W_{b,\ell}=T_{b,\ell}-\tilde{T}_{b,\ell}$.
Finally, the discoveries obtained with our method will be interpreted as highlighting groups of SNPs displaying significant associations with the outcome within certain subgroups determined by specific covariate values.
@@
Background on inference at the level of groups of SNPs.
As discussed in Appendix A2.3.4 and A2.3.4, Sesia et al. (2021) partition the individual SNPs into contiguous groups at seven levels of resolution based on complete linkage hierarchical clustering. The group sizes range from single SNPs to 425-kilobase (kb) wide groups. For the real data analysis we choose the resolution with median group width 81 kb and for the simulation we choose a resolution with median group width 3 kb. In the simulation, for simplicity, we only choose a single variant within each of these groups to be nonzero. Note that the presence of a single nonzero SNP within a group renders the conditional independence hypothesis relative to that group of SNPs to be false as well.
@@
The GWAS data analyses in Sections 3.2 and 3.3 are conducted in R/4.2.0 and use the “bigstatsr” and “bigsnpr” packages (Privé et al., 2018) for handling large-scale genetic data. For data manipulation, transformation and visualization we use the “tidyverse” package (Wickham et al., 2019). Computational details on the construction of knockoff variables based on Hidden Markov Models (HMM) can be found in Sesia et al. (2021). Their software to construct these knockoffs is available at https://msesia.github.io/knockoffgwas/. For reference, the construction of the HMM knockoffs in Sesia et al. (2021) took $\approx$ 4 d using 10 cores and 80 GB of memory to generate knockoffs on chromosome 1 for the UK Biobank data ($\approx$ 1 million haplotype sequences, 600,000 SNPs, and 600,000 IBD (identical-by-descent segments)). In this paper, we run all jobs on a computing cluster and we parallelize when possible. For the simulation in Section 3.2, we submit separate jobs for each amplitude, sparsity, proportion of main effects. Runtime for 100 iterations varies by the simulation parameters and ranges from around 2 hours to 12 hours, with a median runtime of around 5 hours. We run all experiments on the computing cluster with 1 core and 25GB RAM. For the real data analysis in section 3.3, we use 20 CPU cores with a max of 200 GB. We describe the additional computational steps for GWAS data in detail in Section A2. The pre-screening step takes around 2 hours, creating the interactions and running the lasso takes around 5 hours, and the lasso on the screened data takes around 3 hours. We submit different jobs to run the local models within each batch, which take around 5 hours. In total, the entire process takes around 15 hours. Our batch scripts containing additional details available at https://github.com/msesia/i-modelx.
@@
Figure A5: Same general simulation setting as A3, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.
Figure A6: Same simulation setting as  A5 with different levels of sparsity.
@@
Figure A9: Same simulation setting as in A7, however, only $Z_{1}$ is used as a covariate. Group 1 indicates the subgroup $Z_{1}=0$ and group 2 indicates $Z_{1}=1$.